If 2trains 18km apart are traveling towards each other 1 with a velocity of 35m/s & the other moving at 25m/s, how long would it be before the trains pass by each other?


1
Expert's answer
2021-05-05T10:56:02-0400

Velocity of train 1, "v_1=35\space m/s"

Velocity of train 2, "v_2=25\space m/s"

Distance between the trains, "s=18\space km=18\times10^3\space m"

Assuming train 2 to be at rest relative to train 1

Relative velocity of train 1 with respect to train 2, "v_{21}=35-(-25)=60\space m/s"

Time taken by train 1 to reach train 2, "t=\dfrac{s}{v_{21}}=\dfrac{18\times10^3}{60}"

"t=300\space s=5\space min"